OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. Write your answer in set brackets with no spaces, such as {-1.2,2.3}. Round the solution to the nearest tenth. x^2-14x+49=3

OpenStudy (anonymous):

\[(x-7)^2=3\] \[x-7=\pm\sqrt{3}\] \[x=7\pm\sqrt{3}\]

OpenStudy (anonymous):

How would i put that in brackets?